OpenStudy (anonymous):

what is the integral of 7/(t^3+3) from x to 1?

OpenStudy (anonymous):

That's a very messy integral. You sure you typed that in correctly?

OpenStudy (anonymous):

do you mean:\[\int\limits_{x}^{1}2/(t^3+3)dx\] ? ^_^

OpenStudy (anonymous):

That doesn't make the integral any more simple. You could fill up several pages calculating that integral unless I'm missing something obvious.

OpenStudy (anonymous):

lol, calm down, let's just check ? :)

OpenStudy (anonymous):

Yes, i did mean what sstarica said

OpenStudy (anonymous):

except the 2 is a 7

OpenStudy (amistre64):

dx ..... is that a typo.....should be "dt" right?

OpenStudy (anonymous):

LOL yeah, sorry >_<

OpenStudy (anonymous):

yes dt

OpenStudy (amistre64):

would splitting the variables help us? dy/dt = 7/(t^3 +3) dy/7 = dt/(t^3+3) ?? maybe not

OpenStudy (anonymous):

figured it out it is just the original function

OpenStudy (amistre64):

.....thatd be a trick :)

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

how you work that out?

OpenStudy (anonymous):

using the first part of the fundamental thm of calc

OpenStudy (amistre64):

ohhh yeah ...the first part of the thing about that stuff.... yeah.... I was gonna do that but you know.... :)